x power m + y power m = 1.


dy/dx = -x/y.


Find the value of m.


 

Grade:12

3 Answers

Chetan Mandayam Nayakar
11 years ago

mxm-1+mym-1(dy/dx)=0,xm-1=-ym-1dy/dx,dy/dx=-(x/y)m-1=-x/y,m=2

mohan kumar
11 years ago

@chetan Gud job man... U R RITE !!!!!!!!!

amit sharma
11 years ago
mohan my friend its very easy to solve this problem the answer is 2....... as i will, explain you in class...